a projectile is travelling at 30` to the horizontal after 2s of its projection and horizontally after 1s more. Calculate the velocity of projection

Grade:11

1 Answers

Eshan
5 years ago
After 3 seconds, the body is travelling horizontally.

Hence\dfrac{usin\theta}{g}=3
\implies usin\theta=30

Also,v_x(2s)=ucos\theta,v_y(2s)=usin\theta-gt=usin\theta-20

Since it travels at30^{\circ} to horizontal at this time,
\dfrac{v_y}{v_x}=\dfrac{30-20}{ucos\theta}=tan30^{\circ}=\dfrac{1}{\sqrt{3}}
\implies ucos\theta=10\sqrt{3}
\implies u=20\sqrt{3}m/s,\theta=60^{\circ}